In ∆ ABC, AD is the median .If AD produced to F, AC produced E ,angle DCE =
3 /2 angle CDF and angle DAC is equal to 20°,then find angle BDA

Grade:9

1 Answers

sreeda
5 years ago
let angle CDF = x° 
then angle ADC = 180 – x (because its linear pair)
also its given angle DCE = 3/2 angle CDF
                                         = 3/2 x
angle DCE = angle ADC + angle DAC (because exterior angle is equal to sum of interior opposite angles)
3/2 x = (180-x) +20
3/2x + x = 200
5/2x = 200
x = 80 ie angle CDF = 80
angle BDA = angle CDF ( vertically opposite angle)
therefore  angle BDA = 80